# A Criterion for Perfectoid Purity and the Rationality of Thresholds

### v1

@@ 1. Introduction
Let $f\in A$ be such that $(p,f)$ forms a regular sequence. We say that $A/f$ is perfectoid pure if there exists a pure extension $A/f\to B$ to a perfectoid ring. If $A/f$ is perfectoid pure, the perfectoid pure threshold $\operatorname{ppt}(A/f,p)$ with respect to $p$ is defined as
@@
An important open problem is whether the converse direction also holds.
@@
We obtain a partial answer to this question in the following theorem, which completely resolves it when $p=2$.
@@
    then $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$. Moreover, $\mathrm{Proj}(\overline{A/f})$ is an RDP K3 surface of Artin invariant $1$ by [DK03]*Theorem 1.1(vii).
@@ Acknowledgements.
The author wishes to express his gratitude to Teppei Takamastu and Linquan Ma for valuable discussion. The author was supported by JSPS KAKENHI Grant number JP24K16889.
@@ 2.2. Frobenius lifts and perfectoid covers
Let $(A,\mathfrak{m})$ be a regular local ring with a finite Frobenius lift $\phi$, and assume $0\neq p\in\mathfrak{m}$. Let $f\in A$. Then there exist a morphism of $\delta$-rings $A\to B$ to a perfect prism $(B,(d))$ such that:
@@
Set $A^{\prime}:=A[X_{0}]$ and define a ring homomorphism $\phi^{\prime}\colon A^{\prime}\to A^{\prime}$ by $\phi^{\prime}|_{A}=\phi$ and $\phi^{\prime}(X_{0})=X_{0}^{p}$; then $A\to A^{\prime}$ is a morphism of $\delta$-rings. Define
@@
By André’s flatness lemma, there exists a $(p,d)$-faithfully flat morphism of perfect prisms $(B^{\prime},(d))\to(B,(d))$ such that $B/(d)$ contains compatible systems of $p$-power roots of $f$ and $p$. Hence $A\to B/(d)$ is also $p$-completely faithfully flat. By [Bhatt20]*Lemma 5.15, this implies that $A\to B/(d)$ is faithfully flat.
Finally, we verify the third condition. Let $C$ be a perfectoid $A/f$-algebra. By the proof of [Yoshikawa25]*Proposition 2.9, it suffices to show that, after replacing $C$ by a $p$-completely faithfully flat extension, there exists an $A$-algebra homomorphism $B/(d)\to C$. After replacing $C$ by an $p$-completely faithfully flat extension, we may assume that $C$ contains a root of every monic polynomial. We define an $A$-algebra homomorphism $A^{\prime}\to C$ by $X_{0}\mapsto p$. Since $\phi^{\prime}$ is finite, there exists an $A$-algebra homomorphism $B^{\prime}\to C$ such that $d\mapsto 0$. Then the natural map $C\to C^{\prime}:=C\widehat{\otimes}_{B^{\prime}/(d)}B/(d)$ is $p$-completely faithfully flat, and there is an induced $A$-algebra homomorphism $B\to C^{\prime}$, as desired. ∎
@@ 3. Splitting-order sequences
  • If $\Phi_{f,\left(s_{0},\ldots,s_{n-1}\right)}$ is not pure, then $\Phi_{f,\left(s_{0},\ldots,s_{n-1},0\right)}$ is also not pure since $Q_{f,\left(s_{0},\ldots,s_{n-1},0\right)}\simeq Q_{f,\left(s_{0},\ldots,s_{n-1}\right)}$. In particular, the set in Definition 3.6 is non-empty.
  • In the setting of Definition 3.6, the splitting-order sequence is invariant modulo $p^{2}$; indeed, if $f\equiv f^{\prime}\pmod{p^{2}}$, then $\boldsymbol{s}(f)=\boldsymbol{s}(f^{\prime})$ by Proposition 3.2.

### v2

@@ 1. Introduction
Let $f\in A$ be such that $(p,f)$ forms a regular sequence. We say that $A/f$ is perfectoid pure if there exists a pure extension $A/f\to P$ to a perfectoid ring. If $A/f$ is perfectoid pure, the perfectoid pure threshold $\operatorname{ppt}(A/f,p)$ with respect to $p$ is defined as
@@
A natural open problem is whether the converse direction also holds.
@@
We obtain a partial answer to this question in the following theorem, which completely resolves the case $p=2$ and the case where $A/f$ is regular. Moreover, since the splitting–order sequence is invariant modulo $p^{2}$ (see Remark 3.8), we deduce that perfectoid purity and the perfectoid pure threshold are also invariant modulo $p^{2}$.
@@
In particular, if $p=2$ or $A/f$ is regular, then perfectoid purity and $\operatorname{ppt}(A/f,p)$ are invariant modulo $p^{2}$; that is, for any $f^{\prime}\in A$ with $f\equiv f^{\prime}\pmod{p^{2}}$, we have that $A/f$ is perfectoid pure if and only if so is $A/f^{\prime}$, and moreover $\operatorname{ppt}(A/f,p)=\operatorname{ppt}(A/f^{\prime},p)$.
@@
    then $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$. Moreover,
    $\mathrm{Proj}(\overline{k[x_{1},x_{2},x_{3},x_{4}]/f})$
    is an RDP K3 surface of Artin invariant $1$ by [DK03]*Theorem 1.1(vii).
@@ Acknowledgements.
The author wishes to express his gratitude to Teppei Takamastu and Linquan Ma for valuable discussion. He is also grateful to Vignesh Jagathese for helpful comments. The author was supported by JSPS KAKENHI Grant number JP24K16889.
@@ 2.2. Frobenius lifts and perfectoid covers
Let $(A,\mathfrak{m})$ be a regular ring with a finite Frobenius lift $\phi$, and assume $0\neq p\in\mathfrak{m}$. Let $f\in A$. Then there exists a morphism of $\delta$-rings $A\to B$ to a perfect prism $(B,(d))$ such that:
@@
Since $A$ is regular, its Frobenius lift $\phi$ is faithfully flat. Set $A^{\prime}:=A[X_{0}]$ and define a ring endomorphism $\phi^{\prime}\colon A^{\prime}\to A^{\prime}$ by $\phi^{\prime}|_{A}=\phi$ and $\phi^{\prime}(X_{0})=X_{0}^{p}$; then $A\to A^{\prime}$ is a morphism of $\delta$-rings. Define
@@
By André’s flatness lemma, there exists a $(p,d)$-faithfully flat morphism of perfect prisms $(B^{\prime},(d))\to(B,(d))$ such that $B/(d)$ contains compatible systems of $p$-power roots of both $f$ and $p$. Hence $A\to B/(d)$ is also $p$-completely faithfully flat. By [Bhatt20]*Lemma 5.15, this implies that $A\to B/(d)$ is faithfully flat.
We now verify the third condition. Let $C$ be a perfectoid $A/f$-algebra. By the proof of [Yoshikawa25]*Proposition 2.9, it suffices to show that, after replacing $C$ by an $\mathfrak{m}$-completely faithfully flat extension, there exists an $A$-algebra homomorphism $B/(d)\to C$.
After such a base change, we may assume that $C$ contains a root of every monic polynomial and is $\mathfrak{m}$-adically complete. Choose a compatible system $\{p^{1/p^{e}}\}_{e}$ of $p$-power roots of $p$ in $C$. Replacing $A$ by its $\mathfrak{m}$-adic completion, we may assume that $A$ is also $\mathfrak{m}$-adically complete. By further base change, we may assume that $A/\mathfrak{m}$ is perfect.
Define an $A$-algebra homomorphism $A^{\prime}\to C$ by sending $X_{0}\mapsto p$. Since $A$ is $p$-adically complete, by taking lifts of a $p$-basis we can choose elements $x_{1},\ldots,x_{N}\in A$ such that $\{\phi_{*}(x_{1}^{e_{1}}\cdots x_{N}^{e_{N}})\mid 0\leq e_{1},\ldots,e_{N}\leq p-1\}$ forms a basis of $\phi_{*}A$ over $A$.
Let $A_{1}\subseteq\phi_{*}A$ be the $A$-submodule generated by $1,x_{1},\ldots,x_{1}^{p-1}$, then it is $A$-subalgebra. Then there exists a monic polynomial $P\in A[T]$ of degree $p$ such that $A[T]/(P)\simeq A_{1}$ as $A$-algebras. Taking a root $\alpha_{1}$ of $P$ in $C$, we obtain an $A$-algebra homomorphism
$A_{1}\simeq A[T]/(P)\xrightarrow{T\mapsto\alpha_{1}}C.$
Repeating this process, we construct an $A$-algebra homomorphism $\phi_{*}A\to C$.
We now define an $A^{\prime}$-algebra homomorphism by
$\phi^{\prime}_{*}A^{\prime}\simeq\phi_{*}A[X_{0}^{1/p}]\xrightarrow{X_{0}^{1/p}\mapsto p^{1/p}}C.$
Iterating this procedure yields an $A^{\prime}$-algebra homomorphism $B^{\prime}\to C$ sending $d$ to 0.
Finally, the natural map $C\to C^{\prime}:=C\widehat{\otimes}_{B^{\prime}/(d)}B/(d)$ is $p$-completely faithfully flat, and there exists an induced $A$-algebra homomorphism $B\to C^{\prime}$, as desired. ∎
@@ 3. Splitting-order sequences
Remark 3.2.
With the notation as in 3.1, fix $f\in A$. Let $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$.
  • The $\overline{A}$-module structure on $Q_{f,\left(0,l_{1},\ldots,l_{n}\right)}$ induces a natural $\overline{A/f}$-module structure. Indeed, we have
    $f(a_{0},\ldots,a_{n})=(0,\,f^{p}(a_{0}^{p}+a_{1}),\,f^{p^{2}}a_{2},\,\ldots,\,f^{p^{n}}a_{n})=0$
    in $Q_{f,\left(0,l_{1},\ldots,l_{n}\right)}$. In particular, the map $\Phi_{f,\left(0,l_{1},\ldots,l_{n}\right)}$ is an $\overline{A/f}$-module homomorphism.
  • If $\overline{A}$ is $F$-finite, then $Q_{f,\left(l_{1},\ldots,l_{n}\right)}$ is a finite $\overline{A}$-module.
@@
  • If $\Phi_{f,\left(s_{0},\ldots,s_{n-1}\right)}$ is not pure, then $\Phi_{f,\left(s_{0},\ldots,s_{n-1},0\right)}$ is also not pure, since
    $Q_{f,\left(s_{0},\ldots,s_{n-1},0\right)}\simeq Q_{f,\left(s_{0},\ldots,s_{n-1}\right)}.$
    In particular, the set appearing in Definition 3.7 is non-empty.
  • If $(A,\mathfrak{m})$ is Noetherian local, then $\Phi_{f,\left(0,l_{1},\ldots,l_{n}\right)}$ is pure if and only if the induced map
    $\Phi_{f,\left(0,l_{1},\ldots,l_{n}\right)}\colon H^{d}_{\mathfrak{m}}(\overline{A/f})\longrightarrow H^{d}_{\mathfrak{m}}(Q_{f,\left(0,l_{1},\ldots,l_{n}\right)})$
    is injective, where $d:=\dim(\overline{A/f})$. This equivalence holds because $\Phi_{f,\left(0,l_{1},\ldots,l_{n}\right)}$ is an $\overline{A/f}$-module homomorphism by Remark 3.2.
  • In the setting of Definition 3.7, the splitting–order sequence is invariant modulo $p^{2}$. Indeed, if $f\equiv f^{\prime}\pmod{p^{2}}$, then $\boldsymbol{s}(f)=\boldsymbol{s}(f^{\prime})$ by Proposition 3.3.
@@ 4. Perfectoid purity versus splitting-order sequences
In particular, if $p=2$ or $A/f$ is regular, then perfectoid purity and $\operatorname{ppt}(A/f,p)$ are invariant modulo $p^{2}$; that is, for any $f^{\prime}\in A$ with $f\equiv f^{\prime}\pmod{p^{2}}$, we have that $A/f$ is perfectoid pure if and only if so is $A/f^{\prime}$, and moreover $\operatorname{ppt}(A/f,p)=\operatorname{ppt}(A/f^{\prime},p)$.
@@
The final assertion follows immediately from (1), (3), and Remark 3.8.
@@
The theory developed for computing the perfectoid pure threshold gives rise to the following interesting example.
Proposition 4.8.
(cf. Example 7.7) Let $p=2$. Let $(A,\mathfrak{m})$ be an unramified complete regular local ring such that $0\neq p\in\mathfrak{m}$ and $A/\mathfrak{m}$ is perfect. Let $f\in A$ be such that $p,f$ form a regular sequence, and set $R:=A/f$. Assume that $R$ is perfectoid pure with $\operatorname{ppt}(R,p)=0$. Let $A\to A_{\infty}$ be a faithfully flat extension to a perfectoid ring containing compatible systems of $p$-power roots of both $f$ and $p$. If $\overline{R}$ is normal, then
$\bigcap_{n\geq 1}(f^{1/p^{\infty}},p^{1-1/p^{n}})A_{\infty}\;\neq\;(f^{1/p^{\infty}},p)A_{\infty}.$
Proof.
Take a Frobenius lift $\phi$ on $A$ and endow $A$ with the corresponding $\delta$-ring structure. Let $(B,(d))$ be the perfect prism associated to $A_{\infty}$. By replacing $A_{\infty}$ with a $p$-completely faithfully flat extension if necessary, we may assume there exists a morphism of $\delta$-rings $A\to B$ (see the proof of Proposition 2.2).
Since $A_{\infty}$ contains a compatible system of $p$-power roots of $f$, there exists $y\in B$ such that $y-f\in(d)$ and $\phi(y)=y^{p}$. Set $R_{\infty}:=A_{\infty}/f^{1/p^{\infty}}$ and $S:=B/(y^{1/p^{\infty}})$. Then $(S,(d))$ is a perfect prism satisfying $S/(d)\simeq R_{\infty}$, and by the proof of Theorem 4.5, the map $R\to R_{\infty}$ is pure.
Since $\operatorname{ppt}(R,p)=0$ and $p=2$, the splitting–order sequence is $\boldsymbol{s}(f)=(0,1,1,\ldots)$. By Proposition 4.2, each map $\Phi_{S,d,\left(s_{0},\ldots,s_{n}\right)}$ is surjective and
$\operatorname{Ker}(\Phi_{S,d,\left(s_{0},\ldots,s_{n}\right)})=p^{1-1/p^{n}}R_{\infty}.$
Suppose, for contradiction, that
$\bigcap_{n\geq 1}(f^{1/p^{\infty}},p^{1-1/p^{n}})A_{\infty}=(f^{1/p^{\infty}},p)A_{\infty}.$
Then the induced map
[…]
is an isomorphism. By functoriality (Proposition 3.5) and the purity of $R\to R_{\infty}$, the map
[…]
is pure. Since $R$ is complete, purity implies that the map $\Phi_{f,\left(\infty\right)}$ splits as an $\overline{R}$-module homomorphism. Because $Q_{f,\left(s_{0},\ldots,s_{n}\right)}\simeq\overline{W}_{n}(R)$, which commutes with restriction, the map
[…]
also splits as $\overline{R}$-modules by Proposition 3.9.
Let $U$ denote the regular locus of $\operatorname{Spec}\overline{R}$. Then $U$ is smooth since $A/\mathfrak{m}$ is perfect. By [Petrov]*Corollary A.2, since $Q_{U,n}$ is flat over $U$ (see [OTY]*Definition 3.1), there exists an integer $n>0$ such that $U\to Q_{U,n}$ splits. Because $\overline{R}$ is normal, the map $\overline{R}\to Q_{R,n}$ also splits; in particular, $R$ is nn–quasi-FF–split. This contradicts the splitting–order sequence $(0,1,1,\ldots)$.
Hence the assumed equality cannot hold, and we obtain
$\bigcap_{n\geq 1}(f^{1/p^{\infty}},p^{1-1/p^{n}})A_{\infty}\neq(f^{1/p^{\infty}},p)A_{\infty},$
as desired. ∎
@@ 5. A computational method for splitting-order sequences
6. Graded case
In this section, we investigate perfectoid purity for graded rings and its relation to the local case at the homogeneous maximal ideal. We begin by proving that the splitting–order sequence behaves compatibly with localization (Proposition 6.2). We then study graded analogues of Calabi–Yau and Fano type cases (Proposition 6.3 and Theorem 6.4), where we establish periodicity and eventual zeros in the splitting–order sequence, leading to results on perfectoid purity and BCM-regularity.
Notation 6.1.
Let $C$ be a complete divisorial valuation ring with uniformizer $p$ such that $C/(p)$ is $F$-finite. Fix a Frobenius lift $\phi_{C}$ on $C$. Let $A:=C[x_{1},\ldots,x_{N}]$, endowed with a Frobenius lift $\phi$ extending $\phi_{C}$ and satisfying $\phi(x_{i})=x_{i}^{p}$ for $1\leq i\leq N$. We give $A$ a graded structure by setting $A_{0}=C$ and $\deg(x_{i})=\mu_{i}\in\mathbb{Z}_{>0}$ for $1\leq i\leq N$. Let $f\in A$ be a homogeneous element such that $(p,f)$ forms a regular sequence, and put $R:=(A/f)^{\wedge p}$. Set $\mathfrak{m}:=(p,x_{1},\ldots,x_{N})=(p,A_{+})$ and $d:=\dim(\overline{A/f})=N-1$.
Since $\Delta(f)\in A$ is homogeneous of degree $p\deg(f)$, the module $Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ has a natural graded structure, and the map $\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ is graded for all $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $1\leq l_{n}\leq p$. We denote by $\boldsymbol{s}(f)=(s_{0},s_{1},\ldots)$ the splitting–order sequence of $f$.
Proposition 6.2.
With the notation as in 6.1, let $\boldsymbol{s}(f/1)$ denote the splitting–order sequence of $f/1\in A_{\mathfrak{m}}$, where the $\delta$-structure on $A_{\mathfrak{m}}$ is induced by $\phi_{\mathfrak{m}}$. Then $\boldsymbol{s}(f)=\boldsymbol{s}(f/1)$. In particular, if each term of $\boldsymbol{s}(f)$ is at most p1p-1, then $R$ is perfectoid pure and
$\operatorname{ppt}(R,p)=\sum_{i\geq 1}\frac{p-1-s_{i}}{p^{i}}$
Proof.
Take integers $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $1\leq l_{n}\leq p$. Since $Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ is a finite $\overline{A}$-module, the purity of $\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ is equivalent to the splitting of $\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}$. Moreover, as $\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ is a graded $\overline{A}$-module homomorphism, its splitting is equivalent to the splitting of its localization
$(\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)})_{\mathfrak{m}}=\Phi_{A_{\mathfrak{m}},f/1,\left(l_{1},\ldots,l_{n}\right)}$
as a homomorphism of $\overline{A_{\mathfrak{m}}}$-modules by [IY]*Lemma 2.10. By definition, this shows that $\boldsymbol{s}(f)=\boldsymbol{s}(f/1)$, as desired.
Finally, the last assertion follows from Theorem 4.5 together with [IY]*Corollary 4.27. ∎
Proposition 6.3.
With the notation as in 6.1, assume $\deg(f)=\mu_{1}+\cdots+\mu_{N}$. If $s_{r}=0$ for some $r>0$, then $s_{i+r}=s_{i}$ for every $i\geq 0$. In particular, $R$ is perfectoid pure.
Proof.
Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a nonzero socle element. By the assumption $\deg(f)=\sum_{i}\mu_{i}$, the element $\eta$ is homogeneous of degree 0.
Fix $n$ with $s_{n}\leq p-1$. Then there exists $\eta_{n}\in H^{d}_{\mathfrak{m}}(\overline{R})\setminus\{0\}$ such that
$\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\Phi_{f,\left(s_{0},\ldots,s_{n-1},s_{n}+1\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{0},\ldots,s_{n-1},s_{n}+1\right)})\bigr)(\eta)$
$=\bigl(F^{n}_{*}H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\cdot F^{n}_{*}f^{s_{n}}}F^{n}_{*}H^{d}_{\mathfrak{m}}(\overline{A/f^{s_{n}+1}})\xrightarrow{V}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{0},\ldots,s_{n-1},s_{n}+1\right)})\bigr)(\eta_{n}).$
In particular, we have
(6.1) $\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\cdot f^{s_{n-1}}}H^{d}_{\mathfrak{m}}(\overline{A/f^{s_{n-1}+1}})\xrightarrow{\Phi_{f,\left(s_{n-1},s_{n}+1\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{n-1},s_{n}+1\right)})\bigr)(\eta_{n-1})$
$=\bigl(F_{*}H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\cdot F_{*}f^{s_{n}}}F_{*}H^{d}_{\mathfrak{m}}(\overline{A/f^{s_{n}+1}})\xrightarrow{V}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{n-1},s_{n}+1\right)})\bigr)(\eta_{n}).$
Moreover,
(6.2) $\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\Phi_{f,\left(s_{0},\ldots,s_{n},s\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{0},\ldots,s_{n},s\right)})\bigr)(\eta)$
$=\bigl(F^{n}_{*}H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\cdot F^{n}_{*}f^{s_{n}}}F^{n}_{*}H^{d}_{\mathfrak{m}}(\overline{A/f^{s_{n}+1}})\xrightarrow{V\circ\Phi_{f,\left(s_{n},s\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{0},\ldots,s_{n},s\right)})\bigr)(\eta_{n}).$
Since $V$ is injective, the map $\Phi_{f,\left(s_{0},\ldots,s_{n},s\right)}$ is pure if and only if the element
$\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\cdot f^{s_{n}}}H^{d}_{\mathfrak{m}}(\overline{A/f^{s_{n}+1}})\xrightarrow{\Phi_{f,\left(s_{n},s\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{n},s\right)})\bigr)(\eta_{n})$
is nonzero.
Since $s_{r}=0$, we have $s_{1},\ldots,s_{r-1}\leq p-1$ by the definition of $\boldsymbol{s}(f)$. Then $\eta_{r}$ satisfies
$\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\Phi_{f,\left(s_{0},\ldots,s_{r-1},1\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{0},\ldots,s_{r-1},1\right)})\bigr)(\eta)$
$=\bigl(F^{r}_{*}H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{V}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{0},\ldots,s_{r-1},1\right)})\bigr)(\eta_{r}).$
Since both $\Phi_{f,\left(s_{0},\ldots,s_{r-1},1\right)}$ and $V$ are graded, the element $\eta_{r}$ is homogeneous of degree 0. Hence there exists $a\in R^{\times}$ such that $\eta_{0}=a\eta_{r}$. Then we have
$\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\Phi_{f,\left(s_{r},s\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(s_{r},s\right)})\bigr)(\eta_{r})=a\,\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\Phi_{f,\left(0,s\right)}}H^{d}_{\mathfrak{m}}(Q_{f,\left(0,s\right)})\bigr)(\eta),$
and hence $s_{r+1}=s_{1}\leq p-1$. By (6.1), we further have $\eta_{1}=a^{p}\eta_{r+1}$. Repeating this argument inductively gives $\eta_{i}=a^{p^{i}}\eta_{i+r}$ and $s_{i}=s_{i+r}$ for all $i\geq 0$.
Therefore $s_{n}\leq p-1$ for all $n$. Applying Proposition 6.2 and Theorem 4.5, we conclude that $R$ is perfectoid pure. ∎
Theorem 6.4.
With the notation as in 6.1, assume that $m:=\mu_{1}+\cdots+\mu_{N}-\deg(f)>0$. Suppose that $K:=\operatorname{Ker}(F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}))$ has finite length, and set
$t:=\min\{\,l\in\mathbb{Z}\mid K_{l}\neq 0\,\}.$
If there exists $r\geq 1$ such that $p^{r}m<t$ and $s_{r}=0$, then the following hold:
  1. (1)
    $s_{i}=0$ for every $i>r$, and in particular, $R$ is perfectoid pure.
  2. (2)
    If $\operatorname{Spec}(\overline{R})\setminus\{\mathfrak{m}\}$ is strongly $F$-regular, then $R_{\mathfrak{m}}$ is perfectoid BCM-regular.
Proof.
Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a nonzero socle element; then $\eta$ is homogeneous of degree m-m.
For each $n$ with $s_{n}\leq p-1$, take elements $\eta_{n}\in H^{d}_{\mathfrak{m}}(\overline{R})$ as in the proof of Proposition 6.3. Then each $\eta_{n}$ is homogeneous of degree $p^{n}m+s_{n}\deg(f)$. Since $p^{r}m=\deg(\eta_{r})<t$, we have $\eta_{r}\notin K$. Therefore,
$\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\xrightarrow{\ \Phi_{f,\left(0,1\right)}\ }H^{d}_{\mathfrak{m}}(Q_{f,\left(0,1\right)})\simeq F_{*}H^{d}_{\mathfrak{m}}(\overline{R})\bigr)(\eta_{r})=(F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to F_{*}H^{d}_{\mathfrak{m}}(\overline{R}))(\eta_{r})\neq 0.$
By (6.2), this implies $s_{r+1}=0$. Repeating the same argument inductively, we obtain (1).
Next, assume that $\operatorname{Spec}(\overline{R})\setminus\{\mathfrak{m}\}$ is strongly $F$-regular. Then the tight closure of 0 in $H^{d}_{\mathfrak{m}}(\overline{R})$, denoted by $0^{*}$, has finite length. By (1), there exists $n$ such that $\eta_{n}\notin 0^{*}$ and $s_{n}=0$.
Let $R\to P$ be an extension to a perfectoid BCM-algebra $P$, and for each $i\geq 0$ set
$\tau_{i}:=\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\longrightarrow H^{d}_{\mathfrak{m}}(\overline{P})\bigr)(\eta_{i}),\quad\text{where }\eta_{0}:=\eta.$
Since
$\operatorname{Ker}\bigl(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{P})\bigr)\subseteq 0^{*},$
we have $\tau_{n}\neq 0$.
Furthermore, by the functoriality of $\Phi$ (see Proposition 3.5), we obtain
$\Phi_{B,d,\left(s_{0},\ldots,s_{n-1},1\right)}(\tau)=V(F^{n}_{*}(\tau_{n}))\neq 0,$
where $(B,(d))$ is a perfect prism satisfying $B/(d)\simeq P$. Hence $\tau\neq 0$, which implies that the map $R\to P$ is pure. Therefore, $R_{\mathfrak{m}}$ is perfectoid BCM-regular, as desired. ∎